# Rigorous Modeling of Droplet Impact Dynamics on Micro-Structured Surfaces: Nonlocal Theory and SPH Simulation of Pancake Bouncing

### v1

@@ 2.1 Modeling of droplet cohesion
The precise mathematical formulation of $A_{\sigma}$ is of critical importance, as its rigorous definition serves as the foundation for ensuring both the algorithm’s generalizability across diverse scenarios and its reliability in practical applications. Now we will establish a quantitative relationship between the parameter $A_{\sigma}$ and surface tension $\sigma$ through theoretical derivation. Notably, the droplet’s curvature radius $R$ is typically much larger than the cohesive force interaction radius $R_{1}$. This scale separation justifies our model simplification that treats the liquid-gas interface as locally planar within the $R_{1}$ domain (i.e., $R\rightarrow\infty$ approximation). While this planar approximation introduces minor errors in the interaction integrals, the accuracy can be systematically improved by reducing $R_{1}$ through mesh refinement. Therefore, we can determine the relationship between $A_{\sigma}$ and surface tension $\sigma$ by studying the tension on a flat surface. Let $\varepsilon$ denote the thickness of the surface. At any given point $\mathbf{x}_{0}$ on the surface, consider an arbitrary unit tangent vector $\mathbf{t}$ at $\mathbf{x}_{0}$. We use $V_{1}$ to describe the surface particles on one side of the normal plane of $\mathbf{t}$, and $V_{2}$ to denote all particles on the opposite side. Here, $V_{1}$ has a thickness of $\varepsilon$, while both $V_{1}$ and $V_{2}$ are strip-shaped regions of length $L(L\gg{R_{1}})$ (see Fig 2). Indeed, the precise relationship between $A_{\sigma}$ and $\sigma$ can be formally established through the following theorem.
Theorem 1 (Effective surface tension generation)
Consider a fluid domain $\Omega$ with constant density $\rho(\mathbf{x})=\rho_{0}$ for all $\mathbf{x}\in\Omega$. Let $\overline{f}_{\mathrm{att}}$ denote the magnitude of the resultant attractive force along direction $\mathbf{t}$ exerted by the fluid in region $V_{2}$ on the surface of $V_{1}$, defined as
$\overline{f}_{\mathrm{att}}=\int_{V_{1}}\int_{V_{2}}\left[\mathbf{f}(\mathbf{x},\mathbf{y})\cdot\mathbf{t}\right]_{+}\,d\mathbf{y}\,d\mathbf{x},$ (7)
where $\cdot$$_{+}$ denotes the positive part, and $\mathbf{f}$ is the pairwise force given in (5). If the coefficient $A_{\sigma}$ in the (5) takes the form
$A_{\sigma}=\frac{4\sigma}{\rho_{0}\varepsilon}\quad\textrm{such that}\quad\mathbf{f}(\mathbf{x},\mathbf{y})=\frac{4\sigma}{\rho_{0}\varepsilon}\rho(\mathbf{y})\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},{R_{1}}),$ (8)
@@
$\overline{f}_{\mathrm{att}}=\sigma L.$ (9)
This establishes that the cohesive force $\mathbf{f}$ generates an effective fluid interface with surface tension coefficient $\sigma$.
Proof 1
@@
$\overline{f}_{\mathrm{att}}$ $=A_{\sigma}\rho_{0}\int_{V_{1}}\int_{V_{2}}\left[\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},R_{1})\cdot\mathbf{t}\right]_{+}\,d\mathbf{y}\,d\mathbf{x}$
$=A_{\sigma}\rho_{0}\int_{V_{1}}\int_{V_{2}}\left[f(|\mathbf{x}-\mathbf{y}|)\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|}\cdot\mathbf{t}\right]_{+}\,d\mathbf{y}\,d\mathbf{x},$ (10)
where we used the relation $f(|\mathbf{x}|)=\partial_{|\mathbf{x}|}W(\mathbf{x},R_{1})$ and $\nabla_{\mathbf{x}}|\mathbf{x}|=\mathbf{x}/|\mathbf{x}|$.
Let $\theta$ be the angle between $\mathbf{t}$ and $\mathbf{y}-\mathbf{x}$. The integrand simplifies to
$\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|}\cdot\mathbf{t}=-\cos\theta.$
By exploiting the symmetry of $V_{1}$, the volume integral of $\mathbf{x}$ in (1) can be reduced to a line integral via $\mathbf{x}=\mathbf{x}_{0}-s\mathbf{t}$. Furthermore, adopting spherical coordinates $(r,\theta,\phi)$ centered at $\mathbf{x}$ with the transformation
$\mathbf{y}=\mathbf{x}+r\begin{pmatrix}-\cos\theta\\ \sin\theta\cos\phi\\ -\sin\theta\sin\phi\end{pmatrix}\textrm{ and }r=|\mathbf{x}-\mathbf{y}|,$
(1) can be reformulated as
$\overline{f}_{\mathrm{att}}$ $=-A_{\sigma}\rho_{0}L\varepsilon\int_{0}^{R_{1}}\int_{0}^{\pi}\int_{0}^{\arccos(s/R_{1})}\int_{s/\cos\theta}^{R_{1}}f(r)\cos\theta r^{2}\sin\theta drd\theta d\phi ds$
$:=-A_{\sigma}\rho_{0}L\varepsilon G_{f},$ (11)
where $G_{f}$ represents the quadruple integral expression. After careful evaluation of this integral (see Appendix A for detailed derivation), we obtain the simplified form
$G_{f}=\frac{\pi}{3}\int_{0}^{R_{1}}f(r)r^{3}dr.$ (12)
Noting that $f(|\mathbf{x}|)=\partial_{|\mathbf{x}|}W(\mathbf{x},h)$, where $W$ is a kernel function, we can further evaluate this expression to yield
$G_{f}=-\frac{1}{4}.$ (13)
Substituting (13) into (1) gives the intermediate result
$\overline{f}_{\mathrm{att}}=\frac{\varepsilon LA_{\sigma}\rho_{0}}{4}.$ (14)
@@
The physical interpretation of this result follows directly from the definition of surface tension: the interfacial force per unit length between regions $V_{1}$ and $V_{2}$ is exactly $\overline{f}_{\mathrm{att}}/L=\sigma$. This demonstrates conclusively that the pairwise cohesive force $\mathbf{f}(\mathbf{x},\mathbf{y})$ defined in (8) generates an effective surface tension with coefficient $\sigma$ at the interface between the fluid regions. This completes the proof.
Remark 1
Theorem 1 is established under the assumption of zero interface curvature. Nevertheless, this approach maintains reasonable accuracy with a mathematically verifiable relative error below 4.5% for scenarios where the radius of curvature is no less than $R_{1}$. The relative error bound was confirmed through the selection of cubic spline kernel functions and numerical integration of Eq. (1). To avoid unnecessary computational complexity associated with curvature corrections, we retain the original cohesive force formulation (8) without modification.
@@ 2.2 Modeling of substrate contact repulsion
Conventional contact algorithms typically require determining both the substrate’s normal direction and the vertical distance to compute repulsive forces (using Equation (1)). However, this requirement becomes overly restrictive for substrates with complex geometries, such as the one illustrated in Figure 1. The method not only incurs substantial computational overhead but also demonstrates numerical instabilities, especially in concave regions where precise distance definition proves particularly challenging. To overcome these limitations, we propose an alternative approach that computes repulsive forces directly between particle pairs. In fact, this strategy effectively preserves the characteristics of the microcone array. Through the implementation of parallel computing techniques, we maintain manageable computational costs while achieving significantly improved stability and broader applicability across various geometric configurations.
@@
Let $\Gamma$ denote the set of all microcone array points, as illustrated in Figure 3. The repulsive force exerted by a boundary particle at position $\mathbf{y}\in\Gamma$ on a fluid particle at position $\mathbf{x}\in\Omega$ is defined as
$\mathbf{F}(\mathbf{x},\mathbf{y})=-p_{\mathrm{rep}}(\mathbf{x},\mathbf{y})\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},R_{2}),$ (15)
where $R_{2}$ represents the radius of repulsive interaction. The scalar function $p_{\mathrm{rep}}(\mathbf{x},\mathbf{y})$, which we interpret as a pseudo-pressure term, takes the form
$p_{\mathrm{rep}}(\mathbf{x},\mathbf{y})$ $=\tilde{\alpha}\rho(\mathbf{y})\bar{v}_{0}^{2}\left(1-\frac{|\mathbf{x}-\mathbf{y}|}{h}\right).$ (16)
The parameter $\tilde{\alpha}$ characterizes substrate properties and is adjustable, while $\bar{v}_{0}$ denotes the mean initial impact velocity of liquid droplets upon contact with the substrate. In this study, we set $\bar{v}_{0}$ = 0.837 m/s. Consequently, the total repulsive force acting on a fluid particle at position $\mathbf{x}$ is obtained by integrating over all boundary particles as follows
$\mathbf{F}(\mathbf{x})$ $=\int_{\Gamma}\mathbf{F}(\mathbf{x},\mathbf{y})d\mathbf{y}=-\int_{\Gamma}p_{\mathrm{rep}}(\mathbf{x},\mathbf{y})\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},R_{2})d\mathbf{y}.$ (17)
We remark that $p_{\mathrm{rep}}$ shares the same physical dimensions as pressure $p$, justifying its interpretation as the pseudo-pressure at point $\mathbf{x}$ induced by boundary point $\mathbf{y}$.
@@ 2.3 Governing equations with nonlocal interface and boundary interactions
$\frac{\mathrm{d}\rho}{\mathrm{d}t}$ $=-\rho\nabla\cdot\mathbf{u},$ (18)
$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}$ $=-\frac{\nabla p}{\rho}+\frac{\mu}{\rho}\nabla^{2}\mathbf{u}+\mathbf{g}+\frac{4\sigma}{\rho^{2}\varepsilon}\int_{\Omega}\rho(\mathbf{y})\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},R_{1})\,\mathrm{d}\mathbf{y}$ (19)
$\quad-\frac{1}{\rho}\int_{\Gamma}p_{\mathrm{rep}}(\mathbf{x},\mathbf{y})\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},R_{2})\,\mathrm{d}\mathbf{y},$ (20)
$\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}$ $=\mathbf{u},$ (21)
@@
$p=\frac{c^{2}\rho_{0}}{\gamma}\left(\left(\frac{\rho}{\rho_{0}}\right)^{\gamma}-1\right).$ (22)
@@ 3.1 Foundations of SPH Methodology
In the Smoothed Particle Hydrodynamics framework, the kernel approximation (also termed integral approximation) of a scalar field $f(\mathbf{x})$ is given by
$f_{I}(\mathbf{x}):=\int_{\Omega}f(\mathbf{x}^{\prime})W(\mathbf{x}-\mathbf{x}^{\prime},h)\,\mathrm{d}\mathbf{x}^{\prime}\approx f(\mathbf{x}),$ (23)
where $W$ denotes the smoothing kernel function with support radius $h$. Introducing the normalized distance $q=|\mathbf{x}-\mathbf{y}|/(h/2)$, we consider two widely used kernel functions:
1). Cubic Spline Kernel [45]:
$W^{\textrm{cs}}(\mathbf{x}-\mathbf{y},h)=\tilde{W}^{\textrm{cs}}(q)=\begin{cases}\sigma_{1}\left[1-\frac{3}{2}q^{2}\left(1-\frac{q}{2}\right)\right],&0\leq q\leq 1,\\ \sigma_{1}(2-q)^{3}/{4},&1<q\leq 2,\\ 0,&q>2,\end{cases}$ (24)
@@
2). Wendland Quintic Kernel [46]:
$W^{\textrm{wq}}(\mathbf{x}-\mathbf{y},h)=\tilde{W}^{\textrm{wq}}(q)=\begin{cases}\sigma_{2}(1-q/2)^{4}(2q+1),&0\leq q\leq 2,\\ 0,&q>2,\end{cases}$ (25)
with $\sigma_{2}=21/[16\pi(h/2)^{3}]$ for three-dimension. The Wendland Quintic Kernel function can constitute a good choice in terms of computational accuracy and effectiveness, since it provides a higher order of interpolation with a computational cost comparable to the quadratic kernel.
@@
$D^{\beta}f]_{I}(\mathbf{x})=-\int_{\Omega}f(\mathbf{x}^{\prime})D^{\beta}_{\mathbf{x}^{\prime}}W(\mathbf{x}-\mathbf{x}^{\prime},h)\,\mathrm{d}\mathbf{x}^{\prime},$ (26)
@@
$\langle f\rangle_{i}$ $=\sum_{j}\frac{m_{j}}{\rho_{j}}f_{j}W_{ij}^{\textrm{wq}},$ (27)
$\langle\nabla\cdot\mathbf{f}\rangle_{i}$ $=-\sum_{j}\frac{m_{j}}{\rho_{j}}\mathbf{f}_{j}\cdot\nabla_{j}W_{ij}^{\textrm{wq}},$ (28)
where $f_{j}:=f(\mathbf{x}_{j})$, $W_{ij}^{\textrm{wq}}:=W^{\textrm{wq}}(\mathbf{x}_{i}-\mathbf{x}_{j},h)$ and $\nabla_{j}W_{ij}^{\textrm{wq}}:=\nabla_{\mathbf{x}_{j}}W^{\textrm{wq}}(\mathbf{x}_{i}-\mathbf{x}_{j},h)$. Here, $i$ denotes the interpolating particle and $j$ refers to the neighbouring particles within the support. $\frac{m_{j}}{\rho_{j}}$ represents the volume of the particle $j$.
In practical applications, modified versions of Eq. (28) are typically employed rather than their original forms. When a derivative function is multiplied or divided by density, it can be incorporated within the summation operator. Specifically, we consider the following two identities [45]:
$[\rho\nabla\cdot\mathbf{f}](\mathbf{x})$ $=\nabla\cdot(\rho\mathbf{f})(\mathbf{x})-\mathbf{f}(\mathbf{x})\cdot\nabla\rho(\mathbf{x}),$ (29)
$\left[\frac{\nabla f}{\rho}\right](\mathbf{x})$ $=\nabla\left(\frac{f}{\rho}\right)(\mathbf{x})+\frac{f(\mathbf{x})}{\rho^{2}(\mathbf{x})}\nabla\rho(\mathbf{x}).$ (30)
@@
$\langle\rho\nabla\cdot\mathbf{u}\rangle_{i}$ $=-\sum_{j}m_{j}\mathbf{u}_{ij}\cdot\nabla_{i}W_{ij}^{\textrm{wq}},$ (31)
$\left\langle\frac{\nabla p}{\rho}\right\rangle_{i}$ $=\sum_{j}m_{j}\left(\frac{p_{i}}{\rho_{i}^{2}}+\frac{p_{j}}{\rho_{j}^{2}}\right)\nabla_{i}W_{ij}^{\textrm{wq}}.$ (32)
Here, $\mathbf{u}_{ij}:=\mathbf{u}_{i}-\mathbf{u}_{j}$. Another common SPH format is as follows
$\left\langle\frac{\mu}{\rho}\nabla^{2}\mathbf{u}\right\rangle_{i}=\sum_{j}\frac{m_{j}(\mu_{i}+\mu_{j})}{(\rho_{i}+\rho_{j})^{2}/4}\frac{\mathbf{x}_{ij}\cdot\nabla_{i}W_{ij}^{\textrm{wq}}}{|\mathbf{x}_{ij}|^{2}+(0.01h_{ij})^{2}}\mathbf{u}_{ij}.$ (33)

### v2

@@ 2.1 Modeling of droplet cohesion
The precise mathematical formulation of $A_{\sigma}$ is of critical importance, as its rigorous definition serves as the foundation for ensuring both the algorithm’s generalizability across diverse scenarios and its reliability in practical applications. Now we will establish a quantitative relationship between the parameter $A_{\sigma}$ and surface tension $\sigma$ through theoretical derivation. Notably, the droplet’s curvature radius $R$ is typically much larger than the cohesive force interaction radius $R_{\mathrm{att}}$. This scale separation justifies our model simplification that treats the liquid-gas interface as locally planar within the $R_{\mathrm{att}}$ domain (i.e., $R\rightarrow\infty$ approximation). While this planar approximation introduces minor errors in the interaction integrals, the accuracy can be systematically improved by reducing $R_{\mathrm{att}}$ through mesh refinement. Therefore, we can determine the relationship between $A_{\sigma}$ and surface tension $\sigma$ by studying the tension on a flat surface. Let $\varepsilon$ denote the thickness of the surface. At any given point $\boldsymbol{x}_{0}$ on the surface, consider an arbitrary unit tangent vector $\boldsymbol{t}$ at $\boldsymbol{x}_{0}$. We use $V_{1}$ to describe the surface particles on one side of the normal plane of $\boldsymbol{t}$, and $V_{2}$ to denote all particles on the opposite side. Here, $V_{1}$ has a thickness of $\varepsilon$, while both $V_{1}$ and $V_{2}$ are strip-shaped regions of length $L(L\gg{R_{\mathrm{att}}})$ (see Fig 2). Indeed, the precise relationship between $A_{\sigma}$ and $\sigma$ can be formally established through the following theorem.
Theorem 2.2 (Effective surface tension generation).
Consider a fluid domain $\Omega$ with constant density $\rho(\boldsymbol{x})=\rho_{0}$ for all $\boldsymbol{x}\in\Omega$. Let $\overline{f}_{\mathrm{att}}$ denote the magnitude of the resultant attractive force along direction $\boldsymbol{t}$ exerted by the fluid in region $V_{2}$ on the surface of $V_{1}$, defined as
(8) $\overline{f}_{\mathrm{att}}=\int_{V_{1}}\int_{V_{2}}\left[\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})\cdot\boldsymbol{t}\right]_{+}\,d\boldsymbol{y}\,d\boldsymbol{x},$
where $\cdot$$_{+}$ denotes the positive part, and $\boldsymbol{f}$ is the pairwise force given in (6). If the coefficient $A_{\sigma}$ in the (6) takes the form
(9) $A_{\sigma}=\frac{4\sigma}{\rho_{0}\varepsilon}\quad\text{such that}\quad\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})=\frac{4\sigma}{\rho_{0}\varepsilon}\rho(\boldsymbol{y})\nabla_{\boldsymbol{x}}W(\boldsymbol{x}-\boldsymbol{y},{R_{\mathrm{att}}}),$
@@
(10) $\overline{f}_{\mathrm{att}}=\sigma L.$
This establishes that the cohesive force $\boldsymbol{f}$ generates an effective fluid interface with surface tension coefficient $\sigma$.
Proof 2.3.
@@
$\overline{f}_{\mathrm{att}}$ $=A_{\sigma}\rho_{0}\int_{V_{1}}\int_{V_{2}}\left[\nabla_{\boldsymbol{x}}W(\boldsymbol{x}-\boldsymbol{y},R_{\mathrm{att}})\cdot\boldsymbol{t}\right]_{+}\,d\boldsymbol{y}\,d\boldsymbol{x}$
(11) $=A_{\sigma}\rho_{0}\int_{V_{1}}\int_{V_{2}}\left[f(|\boldsymbol{x}-\boldsymbol{y}|)\frac{\boldsymbol{x}-\boldsymbol{y}}{|\boldsymbol{x}-\boldsymbol{y}|}\cdot\boldsymbol{t}\right]_{+}\,d\boldsymbol{y}\,d\boldsymbol{x},$
where we used the relation $f(|\boldsymbol{x}|)=\partial_{|\boldsymbol{x}|}W(\boldsymbol{x},R_{\mathrm{att}})$ and $\nabla_{\boldsymbol{x}}|\boldsymbol{x}|=\boldsymbol{x}/|\boldsymbol{x}|$.
Let $\theta$ be the angle between $\boldsymbol{t}$ and $\boldsymbol{y}-\boldsymbol{x}$. The integrand simplifies to
$\frac{\boldsymbol{x}-\boldsymbol{y}}{|\boldsymbol{x}-\boldsymbol{y}|}\cdot\boldsymbol{t}=-\cos\theta.$
By exploiting the symmetry of $V_{1}$, the volume integral of $\boldsymbol{x}$ in (2.3) can be reduced to a line integral via $\boldsymbol{x}=\boldsymbol{x}_{0}-s\boldsymbol{t}$. Furthermore, adopting spherical coordinates $(r,\theta,\phi)$ centered at $\boldsymbol{x}$ with the transformation
$\boldsymbol{y}=\boldsymbol{x}+r\begin{pmatrix}-\cos\theta\\ \sin\theta\cos\phi\\ -\sin\theta\sin\phi\end{pmatrix}\text{ and }r=|\boldsymbol{x}-\boldsymbol{y}|,$
(2.3) can be reformulated as
$\overline{f}_{\mathrm{att}}$ $=-A_{\sigma}\rho_{0}L\varepsilon\int_{0}^{R_{\mathrm{att}}}\int_{0}^{\pi}\int_{0}^{\arccos(s/R_{\mathrm{att}})}\int_{s/\cos\theta}^{R_{\mathrm{att}}}f(r)\cos\theta r^{2}\sin\theta drd\theta d\phi ds$
(12) $:=-A_{\sigma}\rho_{0}L\varepsilon G_{f},$
where $G_{f}$ represents the quadruple integral expression. By use of Lemma 2.1, we obtain the simplified form
(13) $G_{f}=\frac{\pi}{3}\int_{0}^{R_{\mathrm{att}}}f(r)r^{3}dr.$
Noting that $f(|\boldsymbol{x}|)=\partial_{|\boldsymbol{x}|}W(\boldsymbol{x},R_{\mathrm{att}})$, it follows from Lemma 2.1 that
(14) $G_{f}=-\frac{1}{4}.$
Substituting (14) into (2.3) gives the intermediate result
(15) $\overline{f}_{\mathrm{att}}=\frac{\varepsilon LA_{\sigma}\rho_{0}}{4}.$
@@
The physical interpretation of this result follows directly from the definition of surface tension: the interfacial force per unit length between regions $V_{1}$ and $V_{2}$ is exactly $\overline{f}_{\mathrm{att}}/L=\sigma$. This demonstrates conclusively that the pairwise cohesive force $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})$ defined in (9) generates an effective surface tension with coefficient $\sigma$ at the interface between the fluid regions. This completes the proof.
Remark 2.4.
Theorem 2.2 is established under the assumption of zero interface curvature. Nevertheless, this approach maintains reasonable accuracy with a mathematically verifiable relative error below 4.5% for scenarios where the radius of curvature is no less than $R_{\mathrm{att}}$. The relative error bound was confirmed through the selection of cubic spline kernel functions and numerical integration of Eq. (2.3). To avoid unnecessary computational complexity associated with curvature corrections, we retain the original cohesive force formulation (9) without modification.
@@ 2.2 Modeling of substrate contact repulsion
In conventional contact algorithms, computing repulsive forces (using Equation (1)) involves both the substrate’s normal direction and the vertical distance. However, this requirement becomes overly restrictive for substrates with complex geometries, such as the one illustrated in Figure 1. A major limitation of this method is its inefficiency in handling surfaces with high curvature. Computational performance is constrained by a critical trade-off: achieving accuracy requires fine spatial discretization, which entails significant computational cost, whereas coarse discretization leads to instability due to inaccurate definitions of the surface normal and distance, a challenge particularly acute in concave regions. To overcome these limitations, we propose an alternative approach that computes repulsive forces directly between particle pairs. In fact, this strategy effectively preserves the characteristics of the microcone array. Through the implementation of parallel computing techniques, we maintain manageable computational costs while achieving significantly improved stability and broader applicability across various geometric configurations.
With this in mind, we now introduce the definition of the repulsive force for our model. Consider two particles located at positions $\boldsymbol{x}$ and $\boldsymbol{y}$, with particle $\boldsymbol{x}$ being subjected to the repulsive force field generated by particle $\boldsymbol{y}$. The potential energy of particle $\boldsymbol{x}$ can be approximated by the function
$\Phi(r)=k_{0}W(r,R_{\mathrm{rep}}),$
where $W(r,R_{\mathrm{rep}})$ is the Wendland quintic kernel for three-dimension, $r=|\boldsymbol{x}-\boldsymbol{y}|$ denotes the inter-particle distance and $R_{\mathrm{rep}}$ represents the radius of repulsive interaction. Once the kernel function $W$ is specified, $k_{0}>0$ is a parameter that depends only on the physical properties of the particles at $\boldsymbol{x}$ and $\boldsymbol{y}$. Dimensional analysis shows that the parameter $k_{0}$ has dimensions of $M][L]^{-1}[T]^{-2}$, which corresponds to pressure or stress. The repulsive force exerted on particle $\boldsymbol{x}$ by particle $\boldsymbol{y}$ is then given by the negative gradient of the potential:
$\boldsymbol{F}(\boldsymbol{x},\boldsymbol{y})=-k_{0}\nabla_{\boldsymbol{x}}W(|\boldsymbol{x}-\boldsymbol{y}|,R_{\mathrm{rep}}).$
Let $\Gamma$ denote the set of all microcone array points, as illustrated in Figure 3. The total repulsive force acting on fluid particle $\boldsymbol{x}$ due to all surrounding boundary particles within a domain $\Gamma$ is obtained by integrating the pairwise force over the domain:
$\boldsymbol{F}(\boldsymbol{x})=\int_{\Gamma}\boldsymbol{F}(\boldsymbol{x},\boldsymbol{y})d\boldsymbol{y}=-k_{0}\int_{\Gamma}\nabla_{x}W(\boldsymbol{x}-\boldsymbol{y},R_{\mathrm{rep}})d\boldsymbol{y}.$
While the cohesion parameter $A_{\sigma}$ is directly derived from the macroscopic surface tension $\sigma$, the repulsion parameter $k_{0}$ lacks a direct macroscopic analogue. The parameter $k_{0}$ is a well-defined macroscopic representation of microscopic repulsive forces. In the numerical simulations presented herein, its value is robustly calibrated through controlled experiments. It serves the critical numerical function of preventing unphysical boundary penetration.
@@ 2.3 Governing equations with nonlocal interface and boundary interactions
(16) $\frac{\mathrm{d}\rho}{\mathrm{d}t}$ $=-\rho\nabla\cdot\boldsymbol{u},$
(17) $\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}t}$ $=-\frac{\nabla p}{\rho}+\frac{\mu}{\rho}\nabla^{2}\boldsymbol{u}+\boldsymbol{g}+\frac{4\sigma}{\rho^{2}\varepsilon}\int_{\Omega}\rho(\boldsymbol{y})\nabla_{\boldsymbol{x}}W(\boldsymbol{x}-\boldsymbol{y},R_{\mathrm{att}})\,\mathrm{d}\boldsymbol{y}$
$\quad-\frac{1}{\rho}k_{0}\int_{\Gamma}\nabla_{\boldsymbol{x}}W(\boldsymbol{x}-\boldsymbol{y},R_{\mathrm{rep}})\,\mathrm{d}\boldsymbol{y},$
(18) $\frac{\mathrm{d}\boldsymbol{x}}{\mathrm{d}t}$ $=\boldsymbol{u},$
@@
(19) $p=\frac{c^{2}\rho_{0}}{\gamma}\left(\left(\frac{\rho}{\rho_{0}}\right)^{\gamma}-1\right).$
@@ 3.1 Foundations of SPH Methodology
In the Smoothed Particle Hydrodynamics framework, the kernel approximation (also termed integral approximation) of a scalar field $f(\boldsymbol{x})$ is given by
(20) $f_{I}(\boldsymbol{x}):=\int_{\Omega}f(\boldsymbol{x}^{\prime})W(\boldsymbol{x}-\boldsymbol{x}^{\prime},h)\,\mathrm{d}\boldsymbol{x}^{\prime}\approx f(\boldsymbol{x}),$
where $W$ denotes the smoothing kernel function with support radius $h$. Introducing the normalized distance $q=|\boldsymbol{x}-\boldsymbol{y}|/(h/2)$, we consider two widely used kernel functions and one modified kernel function:
1). Cubic Spline Kernel [monaghan1992smoothed]:
(21) $W^{\text{cs}}(\boldsymbol{x}-\boldsymbol{y},h)=\tilde{W}^{\text{cs}}(q)=\begin{cases}\sigma_{1}\left[1-\frac{3}{2}q^{2}\left(1-\frac{q}{2}\right)\right],&0\leq q\leq 1,\\ \sigma_{1}(2-q)^{3}/{4},&1<q\leq 2,\\ 0,&q>2,\end{cases}$
@@
2). Wendland Quintic Kernel [gomez2010state]:
(22) $W^{\text{wq}}(\boldsymbol{x}-\boldsymbol{y},h)=\tilde{W}^{\text{wq}}(q)=\begin{cases}\sigma_{2}(1-q/2)^{4}(2q+1),&0\leq q\leq 2,\\ 0,&q>2,\end{cases}$
with $\sigma_{2}=21/[16\pi(h/2)^{3}]$ for three-dimensions. The Wendland Quintic Kernel function can constitute a good choice in terms of computational accuracy and effectiveness, since it provides a higher order of interpolation with a computational cost comparable to the quadratic kernel.
3). New Wendland Kernel:
(23) $W^{\text{new}}(\boldsymbol{x}-\boldsymbol{y},h)=\tilde{W}^{\text{new}}(q)=\begin{cases}\sigma_{3}(1-q/2)^{5}(5q/2+1),&0\leq q\leq 2,\\ 0,&q>2,\end{cases}$
with $\sigma_{3}=189/[128\pi(h/2)^{3}]$ for three-dimension. The newly proposed kernel function is a straightforward modification of the Wendland Quintic Kernel, achieved by increasing the polynomial degree, which brings it closer to approximating a repulsive potential. In practice, such a kernel is relatively more effective at capturing the characteristics of short-range forces.
@@
(24) $D^{\beta}f]_{I}(\boldsymbol{x})=-\int_{\Omega}f(\boldsymbol{x}^{\prime})D^{\beta}_{\boldsymbol{x}^{\prime}}W(\boldsymbol{x}-\boldsymbol{x}^{\prime},h)\,\mathrm{d}\boldsymbol{x}^{\prime},$
@@
(25) $\langle f\rangle_{i}$ $=\sum_{j}\frac{m_{j}}{\rho_{j}}f_{j}W_{ij}^{\text{wq}},$
(26) $\langle\nabla\cdot\boldsymbol{f}\rangle_{i}$ $=-\sum_{j}\frac{m_{j}}{\rho_{j}}\boldsymbol{f}_{j}\cdot\nabla_{j}W_{ij}^{\text{wq}},$
where $f_{j}:=f(\boldsymbol{x}_{j})$, $W_{ij}^{\text{wq}}:=W^{\text{wq}}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j},h)$ and $\nabla_{j}W_{ij}^{\text{wq}}:=\nabla_{\boldsymbol{x}_{j}}W^{\text{wq}}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j},h)$. Here, $i$ denotes the interpolating particle and $j$ refers to the neighbouring particles within the support. $\frac{m_{j}}{\rho_{j}}$ represents the volume of the particle $j$.
In practical applications, modified versions of Eq. (26) are typically employed rather than their original forms. When a derivative function is multiplied or divided by density, it can be incorporated within the summation operator. Specifically, we consider the following two identities [monaghan1992smoothed]:
(27) $[\rho\nabla\cdot\boldsymbol{f}](\boldsymbol{x})$ $=\nabla\cdot(\rho\boldsymbol{f})(\boldsymbol{x})-\boldsymbol{f}(\boldsymbol{x})\cdot\nabla\rho(\boldsymbol{x}),$
(28) $\left[\frac{\nabla f}{\rho}\right](\boldsymbol{x})$ $=\nabla\left(\frac{f}{\rho}\right)(\boldsymbol{x})+\frac{f(\boldsymbol{x})}{\rho^{2}(\boldsymbol{x})}\nabla\rho(\boldsymbol{x}).$
@@
(29) $\langle\rho\nabla\cdot\boldsymbol{u}\rangle_{i}$ $=-\sum_{j}m_{j}\boldsymbol{u}_{ij}\cdot\nabla_{i}W_{ij}^{\text{wq}},$
(30) $\left\langle\frac{\nabla p}{\rho}\right\rangle_{i}$ $=\sum_{j}m_{j}\left(\frac{p_{i}}{\rho_{i}^{2}}+\frac{p_{j}}{\rho_{j}^{2}}\right)\nabla_{i}W_{ij}^{\text{wq}}.$
Here, $\boldsymbol{u}_{ij}:=\boldsymbol{u}_{i}-\boldsymbol{u}_{j}$. Another common SPH format is as follows
(31) $\left\langle\frac{\mu}{\rho}\nabla^{2}\boldsymbol{u}\right\rangle_{i}=\sum_{j}\frac{m_{j}(\mu_{i}+\mu_{j})}{(\rho_{i}+\rho_{j})^{2}/4}\frac{\boldsymbol{x}_{ij}\cdot\nabla_{i}W_{ij}^{\text{wq}}}{|\boldsymbol{x}_{ij}|^{2}+(0.01h_{ij})^{2}}\boldsymbol{u}_{ij}.$